4x-3 The function g is defined by g(x) = x+1 Find g(n+3). 8 (n + 3) = 1 xo?
Answer #1

g(x) = ( 4x - 3) / ( x+1)

g(n+3)

plug x = n +3

g(n+3) = ( 4* ( n+3))-3 / ( n+3 + 1)

= ( 4n+9) / ( n + 4)

g(n+3) = ( 4n + 9) / ( n + 4)